OpenStudy (amanafc):

Find the natural abundance of Si-30.

OpenStudy (anonymous):

this is more math than anything... First you need to know how many isotopes there are of silicon, and it's average atomic units (look at periodic table). Then make up a system of equations to solve for it. There's 3 stable silicon isotopes (28, 29, 30) so you will need to have 3 equations. You must be given the percent abundance of at least one of the isotopes to solve because here I can only see 2 equations (numbered down below) set x = percent abundance of si-28 y = percent abundance of si-29 z = percent abundance of si-30 since all of silicon atoms account for 100% of all silicon: x + y + z = 100% = 1 therefore: 1) x = 1 - y - z You also have 2) 28x + 29y + 30z = average atomic mass you can substitute x so that equation becomes: 28 (1 - y - z) + 29y + 30z = average atomic mass See how you have 2 variables here? You can't go on until you know the value of one isotope already or you have given a clue which you can derive the third equation